OpenStudy (anonymous):

1/x+1/y=1/z Solve for y please Thanks

OpenStudy (jamesj):

Notice 1/y = 1/z - 1/x = (x-z)/xz So y = xz/(x-z)

jimthompson5910 (jim_thompson5910):

\[\Large \frac1x+\frac1y=\frac1z\] \[\Large yz+xz=xy \] \[\Large xz=xy-yz\] \[\Large xz=y(x-z)\] \[\Large \frac{xz}{x-z}=y\] \[\Large y=\frac{xz}{x-z}\]

OpenStudy (anonymous):

wow, thanks to both of you.